carl Page 6 of 8 Unit 1 Group Test 7. You are given a function f, a table of values for g and a graph for h. Show all work x g(x) hx) f(x) 2x -2 12 If possible, use the above information to find the following. f(g(h(-2))) a. b. gihf(-1)) f(f(h(5))) d, The weekly cost C of producing x units in a manufacturing process is given by C(x) 600+50x The number of units x produced in t hours is x(t) 40t +10 8. a. Find C(x(t) and simplify. What doesC(x(t)) compute? b. Find the cost of production for 8 hours.
Answer #1

7)

a) f(g(h(-2))) = f(g(4)) = f(0) = 2*0 - 2 = -2

b) f(h(g(4))) = f(h(0)) = f(6) = 2*6 - 2 =10

c) g(h(f(-1))) = g(h(2*(-1) - 2)) = g(h(-4)) = g(2) =10

d) f(f(h(5))) = f(f(-4)) = f((2*(-4)-2)) = f(-10) = 2*(-10) - 2 = -22

8)

a)C(x(t)) = 600 + 50x(t) = 600 + 50(40(t) + 10) = 1100 + 2000t
C(x(t)) computes the cost of producing the units for t hours

b) cost of prouduction for 8 hours = 1100 + 2000*8 = 17,100